OpenStudy (4meisu):

Find k in the expansion (m-2n)^10 = m^10 - 20m^9n + km^8n^2 - .... + 1024n^10

OpenStudy (anonymous):

note that\[(m-2n)^{10}=\sum_{l=0}^{10} \left(\begin{matrix}10 \\ l\end{matrix}\right) m^{10-l} (-2n)^l\]

OpenStudy (anonymous):

Use binomial expansions. \[\left(\begin{matrix}10 \\ 2\end{matrix}\right)(-2)^2 = k\]